Length of brass if the linear expansivity is 1.20×10^–5 kelvin and temperature is 0°c


1
Expert's answer
2021-02-23T10:03:45-0500
"L=L_0(1+\alpha(T-T_0))\\
\alpha=1.2\cdot10^{-5}K^{-1}\\T_0=273.15\ K"